A cube made of lead measures 3 cm on each side and a mass of 310.50 grams. What is its density in g / cm3?

a. 10.35 g/cm3 b. 11.50 g/cm3 c. 34.5 g/cm3 d. 103.5 g/cm3 



1
Expert's answer
2021-04-19T17:09:16-0400

The density is mass divided by volume:


"\rho=\frac mV=\frac{m}{a^3}=\frac{310.50}{3^3}=34.5\text{ g/cm}^3."